OpenStudy (anonymous):

plz tell me the integration of sin(x)/3+sin(2x)

OpenStudy (gorv):

\[\int\limits \frac{ sinx }{ 3} *dx +\int\limits 2sinx*cosx *dx\]

OpenStudy (gorv):

integration of sin = -cos

OpenStudy (gorv):

-cosx/3

OpenStudy (gorv):

sin2x=2sinx*cosx it is an identity

OpenStudy (gorv):

so now we have 2sinx*cosx

OpenStudy (gorv):

let sinx=t differentiate cosx*dx=dt

OpenStudy (gorv):

\[\int\limits2*sinx*cosx*dx=2* \int\limits t*dt\]

OpenStudy (gorv):

arpit can u solve it now

OpenStudy (anonymous):

yes i can solve it thxxx Gorv

OpenStudy (gorv):

welcm:)

OpenStudy (zzr0ck3r):

Notice that the integral of sin(2x) is -(1/2)cos(2x), to see this without subbing just take the derivative of the result, and see how I just did the opposite for the chain rule. (its a good trick)

OpenStudy (zzr0ck3r):

It is easy to undo the chain rule, if only constants are involved

OpenStudy (zzr0ck3r):

@Arpit97 you should give @gorv a medal for helping you understand.